A Bibliography of Publications about the Java Programming Language, 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

27 April 2024 Version 1.36

Abstract

This bibliography records books about the Java Programming Language and related software.

Title word cross-reference

K [Nig22].

-Means [Nig22].

10-year [BPLFRL20].

2019 [APA^{+}20]. 2022 [SNA^{+}23].

3.0 [BM23].

8 [FRD20, KTB20].

abnormality [AAYK20]. Abstraction [DB23, MV20]. according [ORPPG20]. adaptable [HLZ^{+}21]. adaptivity [CBPC23]. Addiction [FLR23]. Adoption [TAV20, FBV22]. Affect [WWW^{+}22]. affected [MNT20]. ahead [MV20]. ahead-of-time [MV20]. aide [CA20]. algorithm [FRD20]. Algorithms [MKNS20, Nig22, GCC20, NGB23]. alternatives [BKP^{+}22]. Analysis [HLX23, KTSS20, SV22, TSBB20, AAYK20, ACSK23, DD20, FHSQ20, FCS20, HLZ^{+}21, HSF^{+}22, LH22, MSDP23]. Analytics [BBB^{+}20, JQZ20]. analyzing [MSB23]. Android [MM22]. Annotations [YBSM21, PVR^{+}20]. ant [GCC20].

Antipatterns [TPBF23]. API [AXR^{+}23, FSZD20, KMSH22, NFH22, ZKX^{+}23, ZWY^{+}20]. APIs [AST23, BLS^{+}23]. application [AAYK20, HSVMB20]. Applications [MM22, RW20, TSBB20, TPBF23, ZBA23, CCRS23, CPV^{+}20, JGSG^{+}21, KAR23, SS23a, TAM^{+}22]. Approach [TSBB20, SAC21, MSDP23]. architectures [PNM^{+}20]. Areas [BM23]. Arithmetic [Dar21]. ARJA [YB20]. ARM [HZN^{+}22]. Art [BM23]. Assessment [MNS^{+}23, TCDT23]. attributes [CAC20]. Automated [MCF^{+}22, TPBF23, YB20, AAYK23, KTB20]. Automatic [KMSH22, MT23, SS23a, SVTTB23, ZKX^{+}23, ZWY^{+}20]. Automatically [AST23]. AWS [Ano20].

Bad [FLR23]. band [DAAZ^{+}20]. Based [BM23, SVDHB23, TPBF23, ASD^{+}23, Agu23, CAC20, CA20, CBPC23, DD20, FHSQ20, FHZ^{+}22, FYL^{+}23, FCS20, HLX23, LFHX23, MSDP23, ZXX23]. behavior [FYL^{+}23]. behavior-based [FYL^{+}23]. BejaGNN [FYL^{+}23]. Benchmarks [AXR^{+}23]. Better [ZXX23]. Beyond [Ram22]. bindings [AS23]. BLAS [FSZD20]. BLASFEO [FSZD20]. Blindspots [BLS^{+}23]. bounded [CK21]. Breaking [FLR23, MNT20]. broken [HFS22]. Bug [ACSK23, AST23]. Build [DBP22]. Bytecode [SVDHB23, FCS20].

C [PVR^{+}20]. calculus [BDGZ23]. Call [HZN^{+}22, ZZ20]. Call-Site [HZN^{+}22]. Capabilities [KMSH22]. Careers [Ano20]. Case [TAV20]. Challenges [BM23]. Changes [WHP^{+}23, MNT20]. Characteristics [HR20, MSB23]. Characterization [SGHM23, RBRB23]. Characterizing [HCL22, YBSM21]. checker [BBG^{+}22]. Checking [WKJ^{+}23]. CIL [FCS20]. classes [SAC21]. classification [ORPPG20]. Clojure [Hic20]. Closer [HZN^{+}22]. Cloud [SV22, NGB23]. Code [AKAS22, BLS^{+}23, HZN^{+}22, MNS^{+}23, MT23, SBBL23, WKJ^{+}23, ZBA23, ACSK23, FHSQ20, FHZ^{+}22, HSF^{+}22, MCF^{+}22, MLBD21, PKB23, PVR^{+}20, TOO^{+}23]. coeffects [BDGZ23]. cognitive [DAAZ^{+}20]. colony [GCC20]. Combining [ASD^{+}23]. Comment [WHP^{+}23]. Commenting [WHP^{+}23]. Comments [WHP^{+}23]. community [BPLFRL20]. Comparative [SV22]. Compilation [HZN^{+}22]. compound [AAYK23]. compression [PKB23]. Concurrency [LMM21]. Conflicts [GMBv20, SGHM23, WWW^{+}22, DBP22]. confusing [AAYK23, TOO^{+}23]. Consistency [HZN^{+}22]. Constraint [TSBB20]. constraints [CFLH^{+}22]. construction [JQZ20]. content [Agu23, CA20]. Context [HLX23, HLZ^{+}21, LH22]. context-sensitive [LH22]. Contextual [DB23]. contracts [HFS22]. control [Ram22]. convex [CK21]. cooperation [CA20]. Coverage [SVDHB23]. Coverage-Based [SVDHB23]. criteria [MSDP23]. critical [CCH^{+}22]. CRNs [DAAZ^{+}20]. cross [CPV^{+}20]. cross-platform [CPV^{+}20]. Cryptographic [AXR^{+}23, ZKX^{+}23]. CvAMoS [DB23].

D [BAP20]. Data [MKNS20, SV22]. dataflow [SS23a]. dataset [KAR23]. Dead [MNS^{+}23, CCRS23]. Debloating [HLX23, SVDHB23, TAM^{+}22]. Debt [ZBA23]. debugging [SIK^{+}21]. decision [MSDP23]. Declarative [C^{+}23, NPZ^{+}20]. decompilation [HSVMB20]. decompiler [HSVMB20]. default [LMM21]. Defect [NXL^{+}22]. Defects [ZWY^{+}20, GCS^{+}20]. definitions [THG20]. Density [WFD23]. Dependencies [SVTTB23]. Dependency [JCA^{+}22, WWW^{+}22]. dependent [CFLH^{+}22]. Deprecation [NFH22]. depth [SBBL23]. Deserialization [SBBL23]. design [BKP^{+}22, PNM^{+}20]. desktop [CCRS23]. despite [HFS22]. Detecting [FHSQ20, MNT20]. Detection [AXR^{+}23, AKAS22, TPBF23, ZKX^{+}23, ZWY^{+}20, AAYK23, ACSK23, FHZ^{+}22, FYL^{+}23, HLZ^{+}21, HSF^{+}22, LFHX23, NBA^{+}21]. Developer [CCH^{+}22, Her21, BPLFRL20]. Developers [MM22, CAC20]. diagrams [SS23a]. Did [MM22]. differ [API^{+}22]. Different [SV22]. directed [FRD20]. Directive [ZWY^{+}20]. diversity [HSVMB20, MLBD21]. do [API^{+}22]. Docker [ZMD21]. Documentation [ZWY^{+}20]. doing [Cob22]. Driven [AST23]. DSL [CBPC23]. DSL-based [CBPC23]. duplex [DAAZ^{+}20]. Dynamic [KTSS20, PKB23, TAM^{+}22].

ecosystem [MPW^{+}21]. effect [CAC20]. Effective [FDD20, TSBB20, LFHX23, TCDT23]. Efficient [SS23b, FDD20, ZZ20]. Elements [FLR23]. Eliminating [MV20]. Elimination [CCH^{+}22, MNS^{+}23]. Empirical [AAYK20, CASA22, MNS^{+}23, HR20, PVR^{+}20, THG20]. employer [Ano20]. end [FBV22]. engine [PKB23]. Entering [BM23]. Environment [SV22, DAAZ^{+}20]. Errors [vO23]. ESLint [TAV20]. Evaluating [KMSH22]. Evaluation [AXR^{+}23, BKP^{+}22, ACSK23, NGB23, PVR^{+}20]. Event [AST23, DB23]. Event-Driven [AST23]. Every [Dar21, Her21]. Evolution [YBSM21, BPLFRL20, CCRS23]. Execution [SBBL23, NPZ^{+}20, PJJM21, SIK^{+}21]. expertise [ORPPG20]. eXPloitation [NBA^{+}21]. Exploits [SBBL23]. exploratory [CCRS23]. expression [Agu23]. extended [NGB23]. extensible [KAR23]. extensive [ACSK23]. extraction [HLZ^{+}21, PJJM21, SS23a]. extractions [HR20]. extractor [Agu23].

fast [LH22]. fault [AAYK20]. fault-prone [AAYK20]. Featherweight [GHK^{+}20]. feature [HLZ^{+}21, LFHX23]. FeynGame [HKL20]. Finding [AST23]. Finite [PJJM21]. Finite-state [PJJM21]. first [WBE20]. Floating [Dar21, ASD^{+}23]. Floating-Point [Dar21, ASD^{+}23]. flow [Ram22]. Forecasting [ZBA23]. Formatting [LFBM23]. formulas [MLBD21]. Framework [WKJ^{+}23, CBPC23, CPV^{+}20, GCC20]. frameworks [FBV22]. frequency [MLBD21]. Friendly [WLC^{+}24]. front [FBV22]. front-end [FBV22]. full [DAAZ^{+}20]. full-duplex [DAAZ^{+}20]. Functional [WFD23, MCF^{+}22, Ram22]. Functions [YAP23, CFLH^{+}22].

Game [MT21]. generate [FRD20]. generators [NPZ^{+}20]. Genetic [YB20]. Git [HHK20]. GitHub [GMBv20]. GMP [AS23]. Go [Fei22, GHK^{+}20]. Google [Fei22]. Graph [MKNS20, FHZ^{+}22, FYL^{+}23, KPK23]. graphical [CPV^{+}20]. Groovy [Kin20]. guided [MCF^{+}22].

HBSniff [HSF^{+}22]. heap [PNM^{+}20]. Heterogeneous [ORPPG20, BDGZ23]. Hibernate [HSF^{+}22]. High [SS23b]. High-Level [SS23b]. Highly [HFS22, AAYK23]. history [Hic20, Kin20]. Hosted [GMBv20]. Hub [ZMD21]. hundreds [MPW^{+}21]. Hybrid [TSBB20].

IBFD [DAAZ^{+}20]. Identification

[MNS^{+}23]. Identifying [MSB23]. IFDS [HLX23]. IFDS-based [HLX23]. illogical [HFS22]. images [ZMD21]. Impact [YBSM21]. imperative [MCF^{+}22, NPZ^{+}20]. implementation [NGB23, PNM^{+}20]. in-band [DAAZ^{+}20]. In-depth [SBBL23]. incremental [LH22]. Independent [KTSS20, WHP^{+}23, JGSG^{+}21]. index [JQZ20]. Information [DB23]. inheritance [BBG^{+}22]. Injection [TSBB20, NBA^{+}21]. Integrated [TSBB20]. integration [THG20]. intelligent [KTB20]. Intensional [C^{+}23]. interactions [BPLFRL20]. Interface [MT21, CPV^{+}20]. Internet [JGSG^{+}21, DAAZ^{+}20]. Interpreter [ZXX23]. Investigation [TOO^{+}23]. Isula [GCC20].

JAVA [WFD23, ASD^{+}23, AXR^{+}23, Agu23, API^{+}22, AAYK20, AAYK23, ACSK23, AS23, Ano20, APA^{+}20, BBG^{+}22, BDGZ23, BKP^{+}22, BPLFRL20, BWTS^{+}23, BLS^{+}23, CCRS23, CAC20, CA20, CBPC23, CK21, CPV^{+}20, C^{+}23, DD20, DAAZ^{+}20, Dar21, Fei22, FRD20, FYL^{+}23, FCS20, FDD20, GCC20, GCS^{+}20, GMBv20, HSVMB20, Her21, HHK20, HR20, HCL22, HSF^{+}22, KAR23, KTB20, KPK23, LFBM23, LH22, MM22, MKNS20, MT23, MV20, MLBD21, MSB23, MT21, Nig22, NGB23, ORPPG20, PJJM21, PNM^{+}20, PVR^{+}20, RBRB23, SBBL23, SS23a, SGHM23, SNA^{+}23, SIK^{+}21, SS23b, SVTTB23, SVDHB23, SAC21, TCDT23, THG20, TPBF23, WKJ^{+}23, WHP^{+}23, WLC^{+}24, YAP23, YBSM21, YB20, ZKX^{+}23, ZWY^{+}20, ZZ20, vO23]. Java-Based [TPBF23]. Java-like [BDGZ23]. Java-type [FDD20]. Java/PCJ [NGB23]. JavaFX [RK20]. JavaScript [AKAS22, AST23, CCH^{+}22, CASA22, FHSQ20, FHZ^{+}22, FBV22, FLR23, HLZ^{+}21, JCA^{+}22, JGSG^{+}21, KTSS20, LFHX23, MSDP23, MNS^{+}23, MNT20, MPW^{+}21, NFH22, NXL^{+}22, NBA^{+}21, PKB23, TAV20, TOO^{+}23, TAM^{+}22, WBE20, ZBA23, ZMD21]. JavaSim [DAAZ^{+}20]. JavaSim-IBFD-CRNs [DAAZ^{+}20]. JEMMA [KAR23]. JGMP [AS23]. JGNN [KPK23]. JGraphT [MKNS20]. JPF [BA23, WKJ^{+}23]. jpf-logic [WKJ^{+}23]. JSAnalyzer [CCH^{+}22]. JSContana [HLZ^{+}21]. JSetL [C^{+}23]. JSON [JQZ20]. JStrong [FHZ^{+}22]. Juliet [ACSK23]. Just [HZN^{+}22, NXL^{+}22]. Just-In-Time [HZN^{+}22, NXL^{+}22]. JVM [BM23]. JVM-Based [BM23].

kernel [DD20]. Key

[BBB^{+}20, HLZ^{+}21, ASD^{+}23]. Kirk [HFS22]. KiWi [BBB^{+}20]. Know [Dar21, Her21]. Kotlin [MM22].

label [ORPPG20]. Landscape [NFH22]. Language [BWTS^{+}23, Fei22, BAP20, Kin20]. Languages [LMM21]. Large [RBRB23, HFS22, HR20, MSB23]. Large-scale [RBRB23]. leak [GCS^{+}20]. Learn [Fei22]. Learned [RK20]. Learning [AST23, SAC21]. Lessons [RK20]. Level [SS23b, ORPPG20]. leveraging [FCS20]. Lexical [CA20]. Libraries [BM23, HCL22]. Library [MKNS20, AS23, CK21, MNT20]. licensing [MPW^{+}21]. light [Agu23]. like [BDGZ23]. limited [SIK^{+}21]. Linters [TAV20, HFS22]. Listen [AST23]. lists [Ano20]. local [AAYK20]. locations [MNT20]. Logic [WKJ^{+}23, WKJ^{+}23]. Look [HZN^{+}22, vO23].

machine [SAC21, DD20, PNM^{+}20]. maintainability [MSB23]. Malicious [AKAS22, FHZ^{+}22, HLZ^{+}21, FHSQ20]. malware [FYL^{+}23]. many [HFS22]. Map [BBB^{+}20]. mapping [HSF^{+}22]. mathematical [MLBD21]. Matrices [FSZD20]. Means [Nig22]. mechanisms [HHK20]. Memory [GCS^{+}20, vO23, PNM^{+}20]. Merge [GMBv20, LFBM23, SGHM23]. meta [HSVMB20]. meta-decompilation [HSVMB20]. Metaverse [BM23]. method [AAYK20, HR20, LFHX23, ZZ20]. methods [CCRS23, HHK20, SAC21]. metrics [SAC21]. MFXSS [LFHX23]. microservice [SS23a]. Middleweight [MT21]. Migrate [MM22]. Migrating [RK20]. migration [JGSG^{+}21]. Mining [NFH22]. mismatches [HFS22]. Misuses [KMSH22, ZKX^{+}23]. ML4Code [KAR23]. Mobile [CCH^{+}22]. model [LFHX23, PJJM21]. Modern [LMM21, THG20]. Multi [YB20, LFHX23, MPW^{+}21]. multi-feature [LFHX23]. multi-licensing [MPW^{+}21]. Multi-Objective [YB20]. Multilevel [DD20]. multiple [CAC20, MSDP23]. Mutating [PVR^{+}20]. mutation [API^{+}22]. My [WWW^{+}22].

names [AAYK23]. native [KPK23]. Nature [GMBv20]. Near [SIK^{+}21]. Near-omniscient [SIK^{+}21]. need [Bie22]. netDFT [WFD23]. network [FHZ^{+}22, FYL^{+}23]. networks [DAAZ^{+}20, KPK23]. Neural [KPK23, FHZ^{+}22, FYL^{+}23, MCF^{+}22]. neural-guided [MCF^{+}22]. NOD4J [SIK^{+}21]. NOde.js [NBA^{+}21]. NodeXP [NBA^{+}21]. Non [CCH^{+}22]. Non-critical [CCH^{+}22]. nonvolatile [PNM^{+}20]. Novel [DAAZ^{+}20]. Novice [BWTS^{+}23]. npm [CASA22].

Obfuscated [AKAS22]. obfuscation [FDD20]. Object [HLX23, HSF^{+}22]. object-relational [HSF^{+}22]. Object-Sensitive [HLX23]. Objective [YB20]. omniscient [SIK^{+}21]. one [MPW^{+}21]. OPA [MSDP23]. Open [GMBv20, THG20]. open-source [THG20]. opportunities [MSB23]. optimization [MV20]. Optimizing [FSZD20]. Ordinal [MSDP23]. Origins [BAP20]. Orthogonal [PNM^{+}20]. OSS [MSB23]. Overflow [BPLFRL20]. overhead [MV20].

Packages [CASA22, MSB23, ZMD21].

Pages [CCH^{+}22]. pairs [AAYK23]. Parallel [MT23, Nig22, NGB23]. parallelization [KTB20]. Party [SVTTB23, HCL22]. Pathfinder [APA^{+}20, SNA^{+}23]. pattern [BKP^{+}22]. Patterns [AST23, TOO^{+}23]. PCJ [NGB23]. pearl [Ram22]. Performance [FSZD20, Nig22, NGB23, TPBF23, ZXX23, TCDT23]. Persistence [Cob22, PNM^{+}20]. persistent [PNM^{+}20]. pipelines [MV20]. Plain [SS23b]. Platform [JGSG^{+}21, KTSS20, CPV^{+}20]. Platform-Independent [KTSS20, JGSG^{+}21]. Point [Dar21, ASD^{+}23]. Pointer [HLX23, LH22]. Polyrun [CK21]. polytopes [CK21]. Practice [TAV20, YBSM21]. Practices [WHP^{+}23]. precise [ZZ20]. predicting [SAC21]. Prediction [NXL^{+}22]. Preservation [LFBM23]. Principal [ZBA23]. priorities [CFLH^{+}22]. Priority [MSDP23]. Program [KMSH22, WWW^{+}22, MCF^{+}22, MV20, PJJM21]. Programmer [Dar21]. Programmers [Fei22, ORPPG20]. Programming [BWTS^{+}23, Cob22, C^{+}23, Fei22, LMM21, SS23b, YB20, BAP20, Kin20]. Programs [YB20, ASD^{+}23, AAYK23, FRD20, MNT20, PVR^{+}20]. Projects [GMBv20, JCA^{+}22, NXL^{+}22, SGHM23, WHP^{+}23, GCS^{+}20, HCL22, MLBD21, SAC21, THG20]. prone [AAYK20]. Properties [WKJ^{+}23]. protection [Bie22]. Python [AAYK23, Ano20, BLS^{+}23, ZXX23, ZMD21].

Quantifying [FLR23].

radio [DAAZ^{+}20]. random [FRD20]. Real [BBB^{+}20, MLBD21]. Real-time [BBB^{+}20]. real-world [MLBD21]. reasoning [ASD^{+}23]. Recommendation [ZWY^{+}20]. Recursive [ZZdSO23, Ram22]. refactoring [KTB20, MSB23]. RegCPython [ZXX23]. Regex [CFLH^{+}22]. Regex-dependent [CFLH^{+}22]. Register [ZXX23]. Register-based [ZXX23]. regular [Agu23]. related [MLBD21]. relational [HSF^{+}22]. Relooper [Ram22]. Remote [SBBL23]. Remote-Code [SBBL23]. Repair [KMSH22, YB20, ZWY^{+}20]. repairs [GCS^{+}20]. Replication [NXL^{+}22, SV22]. representation [FHZ^{+}22]. Research [BM23]. resolution [ZZ20]. resource [GCS^{+}20]. Result [BLS^{+}23]. rich [SS23a]. right [Cob22]. risks [HCL22]. Ruby [ZMD21]. rule [ASD^{+}23]. rule- [ASD^{+}23]. Runtime [WLC^{+}24, CBPC23]. Rust [vO23].

Safe [KTB20, LMM21]. Safe-by-default [LMM21]. Safety [vO23]. sampling [CK21]. Scalable [BBB^{+}20, JQZ20]. scale [HR20, RBRB23]. Security [TSBB20, SS23a]. security-rich [SS23a]. semantic [FHSQ20, FHZ^{+}22]. Semantics [MT21, WWW^{+}22, FCS20]. Semantics-based [FCS20]. Sensitive [HLX23, LH22]. Sequential [MT23]. server [NBA^{+}21, TAM^{+}22]. server-side [NBA^{+}21, TAM^{+}22]. Sets [C^{+}23]. SGX [WLC^{+}24]. SGX-Friendly [WLC^{+}24]. SHARP [LH22]. Should [Dar21, Her21]. side [NBA^{+}21, TAM^{+}22]. similar [AAYK23]. Simplifying [CCH^{+}22]. simulator [DAAZ^{+}20]. Site [HZN^{+}22]. size [SIK^{+}21]. size-limited [SIK^{+}21]. Skills [Ano20]. Slicing [TSBB20]. Small [FSZD20]. smell [HSF^{+}22]. Smells [JCA^{+}22]. SMT [ASD^{+}23]. SMT-based [ASD^{+}23]. social [BPLFRL20]. Software [TPBF23, CAC20, CA20, MSDP23, SAC21, TCDT23]. solid [WFD23]. Solver [MSDP23]. Solving [CFLH^{+}22, TSBB20]. Source [GMBv20, THG20]. Specialization [SVTTB23]. specifications [NPZ^{+}20]. Spork [LFBM23]. spotting [HFS22]. spread [CCRS23]. Sql [Ano20]. Stack [BPLFRL20]. State [BM23, PJJM21, TCDT23]. stateful [JGSG^{+}21]. Static [AXR^{+}23, ACSK23, FCS20, HSF^{+}22]. steady [TCDT23]. still [THG20]. Strategies [SV22]. stream [MV20]. streams [KTB20, RBRB23]. String [YAP23, CFLH^{+}22]. structural [CAC20, JQZ20]. structure [ORPPG20]. Structured [LFBM23, Ram22]. Structures [MKNS20]. Stubbifier [TAM^{+}22]. Study [CASA22, GMBv20, NFH22, NXL^{+}22, SBBL23, SGHM23, TAV20, AAYK20, BPLFRL20, CCRS23, CAC20, CA20, HR20, THG20]. Suboptimal [WHP^{+}23]. Subtyping [ZZdSO23]. Suites [ACSK23]. Support [YAP23]. supporting [BBG^{+}22]. Survey [NFH22]. Swing [RK20]. synthesis [MCF^{+}22].

Taint [KTSS20]. Technical [ZBA23].

Temporal [WKJ^{+}23]. Test

[ACSK23, THG20]. Testing [RW20]. TFA [ZZ20]. Theia [BA23]. their [GCS^{+}20, ORPPG20]. Theory [WFD23]. There [ZKX^{+}23, TCDT23]. Things [Her21, DAAZ^{+}20, JGSG^{+}21]. ThingsMigrate [JGSG^{+}21]. Third [SVTTB23, HCL22]. Third-Party [SVTTB23, HCL22]. Time [HZN^{+}22, NXL^{+}22, BBB^{+}20, MV20]. too [HFS22]. Tool [CCH^{+}22, HSF^{+}22, SIK^{+}21]. Tools [AXR^{+}23, BM23, API^{+}22, ACSK23]. top [Ano20]. Trace [BA23, SIK^{+}21]. Traces [BA23, DD20]. tracking [HHK20]. transducers [CFLH^{+}22]. Transformer [MT23]. translation [FCS20, Ram22]. transpilation [MCF^{+}22]. tree [ORPPG20]. Trivial [CASA22]. Two [Her21]. type [FRD20, FDD20, HFS22]. type-directed [FRD20]. typed [FRD20]. types [HFS22]. typestate [BBG^{+}22].

un- [FBV22]. Understanding [BPLFRL20]. Unifying [NPZ^{+}20]. unit [THG20]. unsound [HFS22]. unstructured [Ram22]. Untriviality [CASA22]. updates [HCL22]. Usage [YBSM21, ZMD21]. usages [HCL22]. Use [BWTS^{+}23]. userspace [DD20]. Using [BA23, C^{+}23, DB23, SAC21, ACSK23, HLZ^{+}21, MSDP23, MCF^{+}22, MV20, SIK^{+}21].

valid [THG20]. value [BBB^{+}20]. variable [AAYK23]. variables [AAYK20, CFLH^{+}22]. verifying [ASD^{+}23]. version [NGB23]. via [FYL^{+}23, YB20]. Viewer [BA23]. Virtual [DD20, PNM^{+}20, ZZ20]. Visual [RW20]. visualization [PJJM21]. Visualize [BA23]. Vulnerabilities [SBBL23]. Vulnerability [AXR^{+}23, TSBB20, LFHX23, NBA^{+}21]. Vulnerable [BLS^{+}23, SAC21].

wanted [Ano20]. WasmView [RW20]. web [Agu23, MSDP23, BM23, CCH^{+}22, FLR23, TSBB20]. web-based [MSDP23]. WebAssembly [RW20]. WebCollectives [Agu23]. well [FRD20]. well-typed [FRD20]. While [RK20]. Wider [YAP23]. wild [DBP22]. Wildcards [Bie22]. Will [WWW^{+}22]. wish [Ano20]. witness [Bie22]. Workshop [APA^{+}20, SNA^{+}23]. world [MLBD21]. wrappers [AS23]. written [SS23a].

XSS [LFHX23].

year [BPLFRL20]. years [WBE20].

ZWT [CPV^{+}20].

References

Aman:2020:ESA

[AAYK20] Hirohisa Aman, Sousuke Amasaki, Tomoyuki Yokogawa, and Minoru Kawahara. Empirical study of abnormality in local variables and its application to fault-prone Java method analysis. Journal of Software: Evolution and Process, 32(4):e2220:1– e2220:??, April 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Aman:2023:ADC

[AAYK23] Hirohisa Aman, Sousuke Amasaki, Tomoyuki Yokogawa, and Minoru Kawahara. An automated detection of confusing variable pairs with highly similar compound names in Java and Python programs. Empirical Software Engineering, 28(5):??, September 2023. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://

link.springer.com/article/
10.1007/s10664-023-10339-
2.

Amankwah:2023:BDJ

[ACSK23] Richard Amankwah, Jinfu Chen, Heping Song, and Patrick Kwaku Kudjo. Bug detection in Java code: an extensive evaluation of static analysis tools using Juliet Test Suites. Software—Practice and Experience, 53(5):1125–1143, May 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Agun:2023:WLR

[Agu23] Hayri Volkan Agun. WebCollectives: a light regular expression based web content extractor in Java. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/
article/pii/S2352711023002650.

Alazab:2022:DOM

[AKAS22] Ammar Alazab, Ansam Khraisat, Moutaz Alazab, and Sarabjot Singh. Detection of obfuscated malicious JavaScript code. Future Internet, 14(8):217, July 22, 2022. CODEN ???? ISSN 1999-5903. URL https://

www.mdpi.com/1999-5903/14/
8/217.

Anonymous:2020:SWS

[Ano20] Anonymous. Skills wanted: Sql, Java, Python, and AWS top employers’ wish lists — [careers]. IEEE Spectrum, 57(1):59, January 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Artho:2020:JPW

[APA^{+}20] Cyrille Artho, Quoc-Sang Phan, Peter Aldous, Alyas Almaawi, Lucas Bang, Lasse Berglund, Tevfik Bultan, Zhenbang Chen, Hayes Converse, Wei Dong, William Eiers, Milos Gligoric, Simon Goldsmith, Lars Grunske, Joshua Hooker, Ismet Burak Kadron, Timo Kehrer, Sarfraz Khurshid, Xuan-Bach D. Le, David Lo, Eric Mercer, Sasa Misailovic, Egor Namakonov, Hoang Lam Nguyen, Yannic Noller, Benjamin Ogles, Rohan Padhye, Pavel Parizek, Corina S. Pasareanu, S. Jacob Powell, Seemanta Saha, Koushik Sen, Elena Sherman, Kyle Storey, Minxing Tang, Willem Visser, Ji Wang, and Hengbiao Yu. The Java Pathfinder Workshop 2019. ACM SIGSOFT Software Engineering Notes, 45 (2):20–22, April 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3385678.3385685.

Amalfitano:2022:HDJ

[API^{+}22] Domenico Amalfitano, Ana C. R. Paiva, Alexis Inquel, Luís Pinto, Anna Rita Fasolino, and René Just. How do Java mutation tools differ? Communications of the ACM, 65(12): 74–89, December 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526099.

Amato:2023:JJB

[AS23] Gianluca Amato and Francesca Scozzari. JGMP: Java bindings and wrappers for the GMP library. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S2352711023001243.

Abbasi:2023:CRS

[ASD^{+}23] Rosa Abbasi, Jonas Schiffl, Eva Darulova, Mattias Ulbrich, and Wolfgang Ahrendt. Combining rule- and SMT-based reasoning for verifying floating-point Java programs in KeY. International Journal on Software Tools for Technology Transfer (STTT), 25(2):185–204, April 2023. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL https://

link.springer.com/article/
10.1007/s10009-022-00691-
x.

Arteca:2023:LHL

[AST23] Ellen Arteca, Max Schäfer, and Frank Tip. Learning how to listen: Automatically finding bug patterns in eventdriven JavaScript APIs. IEEE Transactions on Software Engineering, 49(1):166–184, January 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Afrose:2023:ESV

[AXR^{+}23] Sharmin Afrose, Ya Xiao, Sazzadur Rahaman, Barton P. Miller, and Danfeng Yao. Evaluation of static vulnerability detection tools with Java cryptographic API benchmarks. IEEE Transactions on Software Engineering, 49(2):485–497, February 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Besseling:2023:UTT

[BA23] Johan Besseling and Cyrille Artho. Using Theia trace viewer to visualize JPF traces. ACM SIGSOFT Software Engineering Notes, 48(1):22–26, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573074.3573081.

Bright:2020:ODP

[BAP20] Walter Bright, Andrei Alexandrescu, and Michael Parker. Origins of the D programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):73:1–73:38, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386323.

Basin:2020:KKV

[BBB^{+}20] Dmitry Basin, Edward Bortnikov, Anastasia Braginsky, Guy Golan-Gueta, Eshcar Hillel, Idit Keidar, and Moshe Sulamy. KiWi: a key–value map for scalable real-time analytics. ACM Transactions on Parallel Computing (TOPC), 7(3): 16:1–16:28, August 2020. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3399718.

Bacchiani:2022:JTC

[BBG^{+}22] Lorenzo Bacchiani, Mario Bravetti, Marco Giunti, João Mota, and António Ravara. A Java typestate checker supporting inheritance. Science of Computer Programming, 221(??): ??, September 1, 2022. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0167642322000776.

Bianchini:2023:JLC

[BDGZ23] Riccardo Bianchini, Francesco Dagnino, Paola Giannini, and Elena Zucca. A Java-like calculus with heterogeneous coeffects. Theoretical Computer Science, 971(??):??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0304397523003766.

Bierhoff:2022:WNW

[Bie22] Kevin Bierhoff. Wildcards need witness protection. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):138:1–138:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563301.

Bijlsma:2022:EDP

[BKP^{+}22] Lex A. Bijlsma, Arjan J. F. Kok, Harrie J. M. Passier, Harold J. Pootjes, and Sylvia Stuurman. Evaluation of design pattern alternatives in Java. Software —Practice and Experience, 52 (5):1305–1315, May 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brun:2023:BPJ

[BLS^{+}23] Yuriy Brun, Tian Lin, Jessie Elise Somerville, Elisha M. Myers, and Natalie Ebner. Blindspots in Python and Java APIs result in vulnerable code. ACM Transactions on Software Engineering and Methodology, 32 (3):76:1–76:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3571850.

Bucur:2023:EMJ

[BM23] Vlad Bucur, , and Liviu-Cristian Miclea. Entering the metaverse from the JVM: The state of the art, challenges, and research areas of JVM-based Web 3.0 tools and libraries. Future Internet, 15(9):305, September 07, 2023. CODEN ???? ISSN 1999-5903. URL https://

                                        www.mdpi.com/1999-5903/15/
                                        9/305.

Blanco:2020:USE

[BPLFRL20] Guillermo Blanco, Roi Pérez-López, Florentino Fdez-Riverola, and Anália Maria Garcia Lourenço. Understanding the social evolution of the Java community in Stack Overflow: a 10-year study of developer interactions. Future Generation Computer Systems, 105 (??):446–454, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0167739X19311884.

Brown:2023:NUJ

[BWTS^{+}23] Neil C. C. Brown, Pierre Weill-Tessier, Maksymilian Sekula, Alexandra-Lucia Costache, and Michael Kölling. Novice use of the Java programming language. ACM Transactions on Computing Education, 23(1): 10:1–10:??, March 2023. CO-DEN ???? ISSN 1946-

6226. URL https://dl.acm.
org/doi/10.1145/3551393.

Cristia:2023:DPI

[C^{+}23] Maximiliano Cristiá et al.

Declarative programming with intensional sets in Java using JSetL. The Computer Journal, 66(3):763–784, March 2023. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//academic.oup.com/comjnl/
article/66/3/763/6521749.

Capiluppi:2020:LCC

[CA20] Andrea Capiluppi and Nemitari Ajienka. Lexical content as a cooperation aide: a study based on Java software. The Journal of systems and software, 164(??):??, June 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S016412122030025X.

Capiluppi:2020:EMD

[CAC20] Andrea Capiluppi, Nemitari Ajienka, and Steve Counsell. The effect of multiple developers on structural attributes: a study based on Java software. The Journal of systems and software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S016412122030073X.

Chowdhury:2022:UTP

[CASA22] Md Atique Reza Chowdhury, Rabe Abdalkareem, Emad Shihab, and Bram Adams. On the untriviality of trivial packages: an empirical study of npm JavaScript packages. IEEE Transactions on Software Engineering, 48(8):2695–2708, August 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Carvalho:2023:DBR

[CBPC23] Tiago Carvalho, João Bispo, Pedro Pinto, and João M. P. Cardoso. A DSL-based runtime adaptivity framework for Java. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S2352711023001929.

Chaqfeh:2022:JWD

[CCH^{+}22] Moumena Chaqfeh, Russell Coke, Jacinta Hu, Waleed Hashmi, Lakshmi Subramanian, Talal Rahwan, and Yasir

                                        Zaki. JSAnalyzer: a Web

developer tool for simplifying mobile Web pages through non-critical JavaScript elimination. ACM Transactions on the Web (TWEB), 16(4):17:1– 17:??, November 2022. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3550358.

Caivano:2023:SED

[CCRS23] Danilo Caivano, Pietro Cassieri, Simone Romano, and Giuseppe Scanniello. On the spread and evolution of dead methods in Java desktop applications: an exploratory study. Empirical Software Engineering, 28(3):??, May 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://

link.springer.com/article/
10.1007/s10664-023-10303-
0.

Chen:2022:SSC

[CFLH^{+}22] Taolue Chen, Alejandro Flores-Lamas, Matthew Hague, Zhilei Han, Denghang Hu, Shuanglong Kan, Anthony W. Lin, Philipp Rümmer, and Zhilin Wu. Solving string constraints with regex-dependent functions through transducers with priorities and variables. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):45:1–45:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498707.

Ciomek:2021:PJL

[CK21] Krzysztof Ciomek and Miłosz Kadziński. Polyrun: a Java library for sampling from the bounded convex polytopes. SoftwareX, 13(??):Article 100659, January 2021. CO-DEN ???? ISSN 2352-

7110.  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S2352711021000042.

Cobbs:2022:PPW

[Cob22] Archie L. Cobbs. Persistence programming: Are we doing this right? ACM Queue: Tomorrow’s Computing Today, 20 (1):24–39, January 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526210.

Cirani:2020:ZNC

[CPV^{+}20] Simone Cirani, Marco Picone, Luca Veltri, Luca Zaccomer, and Francesco Zanichelli. ZWT: a new cross-platform graphical interface framework for Java applications. SoftwareX, 12(??): Article 100599, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S2352711020303125.

Darabkh:2020:JIC

[DAAZ^{+}20] Khalid A. Darabkh, Oswa M. Amro, Raed T. Al-Zubi, Haythem Bany Salameh, and Ramzi Saifan. JavaSim-IBFD-CRNs: Novel Java simulator for in-band full-duplex cognitive radio networks over Internet of Things environment. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S1084804520303039.

Darcy:2021:FPA

[Dar21] Joseph D. Darcy. Floatingpoint arithmetic: What every Java programmer should know! Web site 29m37s video.,

2021. URL https://youtu.be/
ajaHQ9S4uTA.

Federico:2023:CEA

[DB23] Gemma Di Federico and Andrea Burattin. CvAMoS — event abstraction using contextual information. Future Internet, 15(3):113, March 18, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.mdpi.
com/1999-5903/15/3/113.

DaSilva:2022:BCW

[DBP22] Léuson Da Silva, Paulo Borba, and Arthur Pires. Build conflicts in the wild. Journal of Software: Evolution and Process, 34(4):e2441:1–e2441:??, April 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Daoud:2020:MAJ

[DD20] Houssem Daoud and Michel Dagenais. Multilevel analysis of the Java Virtual Machine based on kernel and userspace traces. The Journal of systems and software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0164121220300698.

Ferreira:2022:AJF

[FBV22] Fabio Ferreira, Hudson Silva Borges, and Marco Tulio Valente. On the (un-)adoption of JavaScript front-end frameworks. Software—Practice and Experience, 52(4):947–966, April 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ferrara:2020:CJB

[FCS20] Pietro Ferrara, Agostino Cortesi, and Fausto Spoto. From CIL to Java bytecode: Semanticsbased translation for static analysis leveraging. Science of Computer Programming, 191 (??):??, June 1, 2020. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167642320300034.

Foket:2020:EEJ

[FDD20] Christophe Foket, Koen De Bosschere, and Bjorn De Sutter. Effective and efficient Javatype obfuscation. Software— Practice and Experience, 50(2): 136–160, February 2020. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feigenbaum:2022:GJP

[Fei22] Barry Feigenbaum. Go for Java Programmers Learn the Google Go Programming Language. Apress, Berkeley, CA, USA, 2022. ISBN 1-4842-7199-8. xxxv + 582 + 61 + 1 pp. LCCN QA76.73.G63 F45 2022.

Fang:2020:DMJ

[FHSQ20] Yong Fang, Cheng Huang, Yu Su, and Yaoyao Qiu. Detecting malicious JavaScript code based on semantic analysis. Computers & Security, 93(??): Article 101764, June 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167404820300481.

Fang:2022:JMJ

[FHZ^{+}22] Yong Fang, Chaoyi Huang, Minchuan Zeng, Zhiying Zhao, and Cheng Huang. JStrong: Malicious JavaScript detection based on code semantic representation and graph neural network. Computers & Security, 118(??):Article 102715, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0167404822001110.

Fouquet:2023:BBQ

[FLR23] Romain Fouquet, Pierre Laperdrix, and Romain Rouvoy. Breaking bad: Quantifying the addiction of Web elements to JavaScript. ACM Transactions on Internet Technology (TOIT), 23(1):22:1–22:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579846.

Feitosa:2020:TDA

[FRD20] Samuel Feitosa, Rodrigo Ribeiro, and Andre Du Bois. A typedirected algorithm to generate random well-typed Java 8 programs. Science of Computer Programming, 196(??): ??, September 15, 2020. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167642320301039.

Frison:2020:BAB

[FSZD20] Gianluca Frison, Tommaso Sartor, Andrea Zanelli, and Moritz Diehl. The BLAS API of BLASFEO: Optimizing performance for small matrices. ACM Transactions on Mathematical Software, 46(2):15:1– 15:36, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3378671.

Feng:2023:BBB

[FYL^{+}23] Pengbin Feng, Li Yang, Di Lu, Ning Xi, and Jianfeng Ma. BejaGNN: behavior-based Java malware detection via graph neural network. The Journal of Supercomputing, 79(14): 15390–15414, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://

link.springer.com/article/
10.1007/s11227-023-05243-

Gavidia-Calderon:2020:IJF

[GCC20] Carlos Gavidia-Calderon and César Beltrán Castañon. Isula: a Java framework for ant colony algorithms. SoftwareX, 11(??): Article 100400, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/
article/pii/S2352711019300639.

Ghanavati:2020:MRL

[GCS^{+}20] Mohammadreza Ghanavati, Diego Costa, Janos Seboek, David Lo, and Artur Andrzejak. Memory and resource leak defects and their repairs in Java projects. Empirical Software Engineering, 25(1):678–718, January 2020. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://

link.springer.com/article/
10.1007/s10664-019-09731-
8.

Griesemer:2020:FG

[GHK^{+}20] Robert Griesemer, Raymond Hu, Wen Kokke, Julien Lange, Ian Lance Taylor, Bernardo Toninho, Philip Wadler, and Nobuko Yoshida. Featherweight Go. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):149:1–149:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428217.

Ghiotto:2020:NMC

[GMBv20] G. Ghiotto, L. Murta, M. Barros, and A. van der Hoek. On the nature of merge conflicts: A study of 2,731 open source Java projects hosted by GitHub. IEEE Transactions on Software Engineering, 46(8):892– 915, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Huang:2022:CUU

[HCL22] Kaifeng Huang, Bihuan Chen, and Yang Liu. Characterizing usages, updates and risks of third-party libraries in Java projects. Empirical Software Engineering, 27(4):??, July 2022. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://

link.springer.com/article/
10.1007/s10664-022-10131-
8.

Hermans:2021:TTE

[Her21] Felienne Hermans. Two things every Java developer should know. IEEE Software, 38(6): 110–112, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hoeflich:2022:HIK

[HFS22] Joshua Hoeflich, Robert Bruce Findler, and Manuel Serrano. Highly illogical, Kirk: spotting type mismatches in the large despite broken contracts, unsound types, and too many linters. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):142:1–142:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563305.

Higo:2020:TJM

[HHK20] Yoshiki Higo, Shinpei Hayashi, and Shinji Kusumoto. On tracking Java methods with Git mechanisms. The Journal of systems and software, 165(??):??, July 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0164121220300522.

Hickey:2020:HC

[Hic20] Rich Hickey. A history of Clojure. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):71:1– 71:46, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/
3386321.

Harlander:2020:F

[HKL20] R. V. Harlander, S. Y. Klein, and M. Lipp. FeynGame. Computer Physics Communications, 256(??):Article 107465, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/S0010465520302186.

He:2023:IBC

[HLX23] Dongjie He, Jingbo Lu, and Jingling Xue. IFDS-based context debloating for object-sensitive pointer analysis. ACM Transactions on Software Engineering and Methodology, 32(4): 101:1–101:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579641.

Huang:2021:JMJ

[HLZ^{+}21] Yunhua Huang, Tao Li, Lijia Zhang, Beibei Li, and Xiaojie Liu. JSContana: Malicious JavaScript detection using adaptable context analysis and key feature extraction. Computers & Security, 104(??): Article 102218, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167404821000420.

Hora:2020:CME

[HR20] Andre Hora and Romain Robbes. Characteristics of method extractions in Java: a large scale empirical study. Empirical Software Engineering, 25(3):1798–1833, May 2020. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://

link.springer.com/article/
10.1007/s10664-020-09809-
8.

Huang:2022:HSA

[HSF^{+}22] Zijie Huang, Zhiqing Shao, Guisheng Fan, Huiqun Yu, Kang Yang, and Ziyi Zhou. HB-Sniff: a static analysis tool for Java Hibernate object-relational mapping code smell detection. Science of Computer Programming, 217(??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167642322000119.

Harrand:2020:JDD

[HSVMB20] Nicolas Harrand, César Soto-Valero, Martin Monperrus, and Benoit Baudry. Java decompiler diversity and its application to meta-decompilation. The Journal of systems and software, 168(??):??, October 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0164121220301151.

Hartley:2022:JTC

[HZN^{+}22] Tim Hartley, Foivos S. Zakkak, Andy Nisbet, Christos Kotselidis, and Mikel Luján. Justin-time compilation on ARM — a closer look at call-site code consistency. ACM Transactions on Architecture and Code Optimization, 19(4):54:1– 54:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL

https://dl.acm.org/doi/10.
1145/3546568.

Jafari:2022:DSJ

[JCA^{+}22] Abbas Javan Jafari, Diego Elias Costa, Rabe Abdalkareem, Emad Shihab, and Nikolaos Tsantalis. Dependency smells in JavaScript projects. IEEE Transactions on Software Engineering, 48(10):3790–3807, October 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Jung:2021:TPI

[JGSG^{+}21] Kumseok Jung, Julien Gascon-Samson, Shivanshu Goyal, Armin Rezaiean-Asel, and Karthik Pattabiraman. ThingsMigrate: Platform-independent migration of stateful JavaScript Internet of Things applications. Software—Practice and Experience, 51(1):117–155, January 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jiang:2020:SSI

[JQZ20] Lin Jiang, Junqiao Qiu, and Zhijia Zhao. Scalable structural index construction for JSON analytics. Proceedings of the VLDB Endowment, 14(4):694– 707, December 2020. CODEN ???? ISSN 2150-8097. URL

                                        https://dl.acm.org/doi/10.
                                        14778/3436905.3436926.

Karmakar:2023:JEJ

[KAR23] Anjan Karmakar, Miltiadis Allamanis, and Romain Robbes. JEMMA: an extensible Java dataset for ML4Code applications. Empirical Software Engineering, 28(2):??, March 2023. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://

link.springer.com/article/
10.1007/s10664-022-10275-
7.

King:2020:HGP

[Kin20]  Paul King.  A history of

the Groovy programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):76:1– 76:53, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/
3386326.

Kechagia:2022:EAP

[KMSH22] Maria Kechagia, Sergey Mechtaev, Federica Sarro, and Mark Harman. Evaluating automatic program repair capabilities to repair API misuses. IEEE Transactions on Software Engineering, 48(7):2658–2679, July 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Krasanakis:2023:JGN

[KPK23] Emmanouil Krasanakis, Symeon Papadopoulos, and Ioannis Kompatsiaris. JGNN: Graph Neural Networks on native Java. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://www.

sciencedirect.com/science/
article/pii/S2352711023001553.

Khatchadourian:2020:SAR

[KTB20] Raffi Khatchadourian, Yiming Tang, and Mehdi Bagherzadeh. Safe automated refactoring for intelligent parallelization of Java 8 streams. Science of Computer Programming, 195(??):

??, September 1, 2020. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167642320300861.

Karim:2020:PID

[KTSS20] R. Karim, F. Tip, A. Soch rková, and K. Sen. Platformindependent dynamic taint analysis for JavaScript. IEEE Transactions on Software Engineering, 46(12):1364–1379, December 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Larsen:2023:PSS

[LFBM23] Simon Larsén, Jean-Rémy Falleri, Benoit Baudry, and Martin Monperrus. Spork: Structured merge for Java with formatting preservation. IEEE Transactions on Software Engineering, 49(1):64–83, January 2023. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Liu:2023:MEX

[LFHX23] Zhonglin Liu, Yong Fang, Cheng Huang, and Yijia Xu. MFXSS: an effective XSS vulnerability detection method in JavaScript based on multi-feature model. Computers & Security, 124 (??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167404822004072.

Liu:2022:SFI

[LH22] Bozhen Liu and Jeff Huang.

SHARP: fast incremental contextsensitive pointer analysis for Java. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):88:1– 88:28, April 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527332.

Liu:2021:SDC

[LMM21] Lun Liu, Todd Millstein, and Madanlal Musuvathi. Safe-by-default concurrency for modern programming languages. ACM Transactions on Programming Languages and Systems, 43(3): 10:1–10:50, September 2021. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3462206.

Mariano:2022:ATI

[MCF^{+}22] Benjamin Mariano, Yanju Chen, Yu Feng, Greg Durrett, and Isil Dillig. Automated transpilation of imperative to functional code using neural-guided program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):71:1–71:27, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527315.

Michail:2020:JJL

[MKNS20] Dimitrios Michail, Joris Kinable, Barak Naveh, and John V.

Sichi. JGraphT — a Java library for graph data structures and algorithms. ACM Transactions on Mathematical Software, 46(2): 16:1–16:29, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381449.

Moseler:2021:DFC

[MLBD21] Oliver Moseler, Felix Lemmer, Sebastian Baltes, and Stephan Diehl. On the diversity and frequency of code related to mathematical formulas in realworld Java projects. The Journal of systems and software, 172 (??):??, February 2021. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0164121220302533.

Martinez:2022:WDD

[MM22] Matias Martinez and Bruno Gois Mateus. Why did developers migrate Android applications from Java to Kotlin? IEEE Transactions on Software Engineering, 48(11):4521–4534, November 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Malavolta:2023:JDC

[MNS^{+}23] Ivano Malavolta, Kishan Nirghin, Gian Luca Scoccia, Simone Romano, Salvatore Lombardi, Giuseppe Scanniello, and Patricia Lago. JavaScript dead code identification, elimination, and empirical assessment. IEEE Transactions on Software Engineering, 49(7):3692–3714, July 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Moller:2020:DLJ

[MNT20] Anders Møller, Benjamin Barslev Nielsen, and Martin Toldam Torp. Detecting locations in JavaScript programs affected by breaking library changes. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA): 187:1–187:25, November 2020.

URL https://dl.acm.org/
doi/10.1145/3428255.

Moraes:2021:OHM

[MPW^{+}21] João Pedro Moraes, Ivanilton Polato, Igor Wiese, Filipe Saraiva, and Gustavo Pinto. From one to hundreds: multilicensing in the JavaScript ecosystem. Empirical Software Engineering, 26(3):??, May 2021. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL http://

link.springer.com/article/
10.1007/s10664-020-09936-
2.

Mumtaz:2023:IRO

[MSB23] Haris Mumtaz, Paramvir Singh, and Kelly Blincoe. Identifying refactoring opportunities for large packages by analyzing maintainability characteristics in Java OSS. The Journal of systems and software, 202(??):??, August 2023. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0164121223001127.

Mahmoudi:2023:OSW

[MSDP23] Amin Mahmoudi, Mahsa Sadeghi, Xiaopeng Deng, and Pengcheng Pan. OPA Solver: a webbased software for Ordinal Priority Approach in multiple criteria decision analysis using JavaScript. SoftwareX, 24 (??):??, December 2023. CO-DEN ???? ISSN 2352-

                                        7110.  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S235271102300242X.

Murawski:2021:GSI

[MT21] Andrzej S. Murawski and Nikos Tzevelekos. Game semantics for interface middleweight Java. Journal of the ACM, 68(1): 4:1–4:51, February 2021. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3428676.

Midolo:2023:ATS

[MT23] Alessandro Midolo, , and Emiliano Tramontana. An automatic transformer from sequential to parallel Java code. Future Internet, 15(9):306, September 08, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.mdpi.
                                        com/1999-5903/15/9/306.

Moller:2020:EAO

[MV20] Anders Møller and Oskar Haarklou Veileborg. Eliminating abstraction overhead of Java stream pipelines using ahead-of-time program optimization. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):168:1–168:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428236.

Ntantogian:2021:NNJ

[NBA^{+}21] Christoforos Ntantogian, Panagiotis Bountakas, Dimitris Antonaropoulos, Constantinos Patsakis, and Christos Xenakis. NodeXP: NOde.js serverside JavaScript injection vulnerability DEtection and eX-Ploitation. Journal of Information Security and Applications (JISA), 58(??):??, May 2021. CODEN ???? ISSN 2214-2126. URL http://www.

sciencedirect.com/science/
article/pii/S221421262100003X.

Nascimento:2022:JAD

[NFH22] Romulo Nascimento, Eduardo Figueiredo, and Andre Hora. JavaScript API deprecation landscape: a survey and mining study. IEEE Software, 39(3): 96–105, May/June 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nowicki:2023:PEJ

[NGB23] Marek Nowicki, Lukasz Górski, and Piotr Bała. Performance evaluation of Java/PCJ implementation of parallel algorithms on the cloud (extended version). Concurrency and Computation: Practice and Experience, 35(15): e6536:1–e6536:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nigro:2022:PPK

[Nig22] Libero Nigro. Performance of parallel K-means algorithms in Java. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/
1999-4893/15/4/117.

Nie:2020:UEI

[NPZ^{+}20] Pengyu Nie, Marinela Parovic, Zhiqiang Zang, Sarfraz Khurshid, Aleksandar Milicevic, and Milos Gligoric. Unifying execution of imperative generators and declarative specifications. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):217:1–217:26, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428285.

Ni:2022:JTD

[NXL^{+}22] Chao Ni, Xin Xia, David Lo, Xiaohu Yang, and Ahmed E. Hassan. Just-in-time defect prediction on JavaScript projects: a replication study. ACM Transactions on Software Engineering and Methodology, 31(4): 76:1–76:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3508479.

Ortin:2020:HTS

[ORPPG20] Francisco Ortin, Oscar Rodriguez-Prieto, Nicolas Pascual, and Miguel Garcia. Heterogeneous tree structure classification to label Java programmers according to their expertise level. Future Generation Computer Systems, 105(??):380–394, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0167739X1931516X.

P:2021:FSM

[PJJM21] Jevitha K. P., Swaminathan Jayaraman, Bharat Jayaraman, and Sethumadhavan M. Finitestate model extraction and visualization from Java program execution. Software—Practice and Experience, 51(2):409– 437, February 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Park:2023:DCC

[PKB23] Hyukwoo Park, Seonghyun Kim, and Boram Bae. Dynamic code compression for JavaScript engine. Software— Practice and Experience, 53(5): 1196–1217, May 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perez:2020:OPN

[PNM^{+}20] Taciano D. Perez, Marcelo V. Neves, Diego Medaglia, Pedro H. G. Monteiro, and César A. F. De Rose. Orthogonal persistence in nonvolatile memory architectures: a persistent heap design and its implementation for a Java Virtual Machine. Software—Practice and Experience, 50(4):368–387, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pinheiro:2020:MCA

[PVR^{+}20] Pedro Pinheiro, José Carlos Viana, Márcio Ribeiro, Leo Fernandes, Fabiano Ferrari, Rohit Gheyi, and Baldoino Fonseca. Mutating code annotations: an empirical evaluation on Java and C programs. Science of Computer Programming, 191(??):??, June 1, 2020. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167642320300290.

Ramsey:2022:BRR

[Ram22]  Norman Ramsey.  Beyond

relooper: recursive translation of unstructured control flow to structured control flow (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 90:1–90:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547621.

Rosales:2023:LSC

[RBRB23] Eduardo Rosales, Matteo Basso, Andrea Rosà, and Walter Binder. Large-scale characterization of Java streams. Software —Practice and Experience, 53 (9):1763–1792, September 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robillard:2020:LLW

[RK20] M. P. Robillard and K. Kutschera.

Lessons learned while migrating from Swing to JavaFX. IEEE Software, 37(3):78–85, May/ June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Romano:2020:WVT

[RW20] A. Romano and W. Wang.

WasmView: Visual testing for WebAssembly applications. In 2020 IEEE/ACM 42nd International Conference on Software Engineering: Companion Proceedings (ICSE-Companion), pages 13–16. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020.

URL https://ieeexplore.
ieee.org/document/9270402.

Sultana:2021:USM

[SAC21] Kazi Zakia Sultana, Vaibhav Anu, and Tai-Yin Chong. Using software metrics for predicting vulnerable classes and methods in Java projects: a machine learning approach. Journal of Software: Evolution and Process, 33(3):e2303:1–e2303:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sayar:2023:DSJ

[SBBL23] Imen Sayar, Alexandre Bartel, Eric Bodden, and Yves Le Traon. An in-depth study of Java deserialization remotecode execution exploits and vulnerabilities. ACM Transactions on Software Engineering and Methodology, 32(1):25:1– 25:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3554732.

Shen:2023:CSM

[SGHM23] Bowen Shen, Muhammad Ali Gulzar, Fei He, and Na Meng. A characterization study of merge conflicts in Java projects. ACM Transactions on Software Engineering and Methodology, 32(2): 40:1–40:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546944.

Shimari:2021:NNO

[SIK^{+}21] Kazumasa Shimari, Takashi Ishio, Tetsuya Kanda, Naoto Ishida, and Katsuro Inoue. NOD4J: Near-omniscient debugging tool for Java using sizelimited execution trace. Science of Computer Programming, 206(??):??, June 1, 2021. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S016764232100023X.

Sherman:2023:JPW

[SNA^{+}23] Elena Sherman, Yannic Noller, Cyrille Artho, Franck van Breugel, Anto Nanah Ji, John Kellerman, Parssa Khazra, Filip Kliber, Gaurang Kudale, Pavel Parízek, Corina S. Pasareanu, Ron Pressler, Matt Walker, Hongru Wang, and Qiuchen Yan. The Java Pathfinder Workshop 2022. ACM SIG-SOFT Software Engineering Notes, 48(1):19–21, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573074.3573080.

Schneider:2023:AES

[SS23a] Simon Schneider and Riccardo Scandariato. Automatic extraction of security-rich dataflow diagrams for microservice applications written in Java. The Journal of systems and software, 202(??):??, August 2023. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0164121223001176.

Silva:2023:EHL

[SS23b] Rui S. Silva and João L. Sobral. Efficient high-level programming in plain Java. International Journal of Parallel Programming, 51(1):22– 42, February 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL https://

link.springer.com/article/
10.1007/s10766-022-00747-
0.

Sasikumar:2022:CAD

[SV22] K. Sasikumar and B. Vijayakumar. Comparative analysis of different data replication strategies in cloud environment. International Journal of Image and Graphics (IJIG), 22 (05):??, October 2022. ISSN 0219-4678. URL https:/

/www.worldscientific.com/
doi/10.1142/S0219467822500425.

Soto-Valero:2023:CBD

[SVDHB23] César Soto-Valero, Thomas Durieux, Nicolas Harrand, and Benoit Baudry. Coverage-based debloating for Java bytecode. ACM Transactions on Software Engineering and Methodology, 32(2):38:1–38:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546948.

Soto-Valero:2023:AST

[SVTTB23] César Soto-Valero, Deepika Tiwari, Tim Toady, and Benoit Baudry. Automatic specialization of third-party Java dependencies. IEEE Transactions on Software Engineering, 49(11):5027–5045, November 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Turcotte:2022:SDD

[TAM^{+}22] Alexi Turcotte, Ellen Arteca, Ashish Mishra, Saba Alimadadi, and Frank Tip. Stubbifier: debloating dynamic serverside JavaScript applications. Empirical Software Engineering, 27(7):??, December 2022. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://

link.springer.com/article/
10.1007/s10664-022-10195-
6.

Tomasdottir:2020:AJL

[TAV20] K. F. Tómasdóttir, M. Aniche, and A. Van Deursen. The adoption of JavaScript linters in practice: A case study on ESLint. IEEE Transactions on Software Engineering, 46(8):863–891, 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Traini:2023:TEA

[TCDT23] Luca Traini, Vittorio Cortellessa, Daniele Di Pompeo, and Michele Tucci. Towards effective assessment of steady state performance in Java software: are we there yet? Empirical Software Engineering, 28(1):??, January 2023. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL https://

link.springer.com/article/

10.1007/s10664-022-10247-
x.

Trautsch:2020:UIT

[THG20] Fabian Trautsch, Steffen Herbold, and Jens Grabowski. Are unit and integration test definitions still valid for modern Java projects? An empirical study on open-source projects. The Journal of systems and software, 159(??):??, January 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0164121219301955.

Torres:2023:ICC

[TOO^{+}23] Adriano Torres, Caio Oliveira, Márcio Okimoto, Diego Marcílio, Pedro Queiroga, Fernando Castor, Rodrigo Bonifácio, Edna Dias Canedo, Márcio Ribeiro, and Eduardo Monteiro. An investigation of confusing code patterns in JavaScript. The Journal of systems and software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0164121223001267.

Trubiani:2023:ADS

[TPBF23] Catia Trubiani, Riccardo Pinciroli, Andrea Biaggi, and Francesca Arcelli Fontana. Automated detection of software performance antipatterns in Java-based applications. IEEE Transactions on Software Engineering, 49(4):2873–2891, April 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Thome:2020:IAE

[TSBB20] J. Thomé, L. K. Shar, D. Bianculli, and L. Briand. An integrated approach for effective injection vulnerability analysis of Web applications through security slicing and hybrid constraint solving. IEEE Transactions on Software Engineering, 46(2):163–195, February 2020. CODEN IESEDJ. ISSN 2326-3881.

vanOorschot:2023:MEMb

[vO23] Paul C. van Oorschot. Memory errors and memory safety: a look at Java and Rust. IEEE Security & Privacy, 21(3):62–68, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Wirfs-Brock:2020:JFY

[WBE20] Allen Wirfs-Brock and Brendan Eich. JavaScript: the first 20 years. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):77:1–77:189, June

2020. URL https://dl.acm.
org/doi/abs/10.1145/3386327.

Wijaya:2023:NJD

[WFD23] Agung Danu Wijaya, Dedy Farhamsa, and Darmawati Darwis. netDFT: JAVA Density Functional Theory for solid. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-

7110.  URL http://www.
sciencedirect.com/science/
article/pii/S2352711023001413.

Wang:2023:SCJ

[WHP^{+}23] Chao Wang, Hao He, Uma Pal, Darko Marinov, and Minghui Zhou. Suboptimal comments in Java projects: From independent comment changes to commenting practices. ACM Transactions on Software Engineering and Methodology, 32(2): 45:1–45:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546949.

Walker:2023:JLF

[WKJ^{+}23] Matt Walker, Parssa Khazra, Anto Nanah Ji, Hongru Wang, and Franck van Breugel. jpflogic: a framework for checking temporal logic properties of Java code. ACM SIG-SOFT Software Engineering Notes, 48(1):32–36, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573074.3573083.

Wu:2024:TSF

[WLC^{+}24] Mingyu Wu, Zhe Li, Haibo Chen, Binyu Zang, Shaojun Wang, Lei Yu, Sanhong Li, and Haitao Song. Toward an SGXfriendly Java runtime. IEEE Transactions on Computers, 73 (1):44–57, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:WDC

[WWW^{+}22] Ying Wang, Rongxin Wu, Chao Wang, Ming Wen, Yepang Liu, Shing-Chi Cheung, Hai Yu, Chang Xu, and Zhiliang Zhu. Will dependency conflicts affect my program’s semantics? IEEE Transactions on Software Engineering, 48(7):2295–2316, July 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yan:2023:TWS

[YAP23] Qiuchen Yan, Cyrille Artho, and Pavel Parízek. Towards wider support for Java string functions. ACM SIG-SOFT Software Engineering Notes, 48(1):37, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573074.3573084.

Yuan:2020:AAR

[YB20] Y. Yuan and W. Banzhaf.

ARJA: Automated repair of Java programs via multiobjective genetic programming. IEEE Transactions on Software Engineering, 46(10):1040– 1067, October 2020. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yu:2021:CUE

[YBSM21] Zhongxing Yu, Chenggang Bai, Lionel Seinturier, and Martin Monperrus. Characterizing the usage, evolution and impact of Java annotations in practice.

IEEE Transactions on Software Engineering, 47(5):969– 986, 2021. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zozas:2023:FPC

[ZBA23] Ioannis Zozas, Stamatia Bibi, and Apostolos Ampatzoglou. Forecasting the principal of code technical debt in JavaScript applications. IEEE Transactions on Software Engineering, 49(4): 2498–2512, April 2023. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zhang:2023:ADJ

[ZKX^{+}23] Ying Zhang, Md Mahir Asef Kabir, Ya Xiao, Danfeng Yao, and Na Meng. Automatic detection of Java cryptographic API misuses: Are we there yet? IEEE Transactions on Software Engineering, 49(1): 288–303, January 2023. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zerouali:2021:UJP

[ZMD21] Ahmed Zerouali, Tom Mens, and Coen De Roover. On the usage of javascript, Python and Ruby packages in Docker Hub images. Science of Computer Programming, 207(??): ??, July 1, 2021. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0167642321000460.

Zhou:2020:ADR

[ZWY^{+}20] Y. Zhou, C. Wang, X. Yan, T. Chen, S. Panichella, and H. Gall. Automatic detection and repair recommendation of directive defects in Java API documentation. IEEE Transactions on Software Engineering, 46(9):1004–1023, September 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zhang:2023:RRB

[ZXX23] Qiang Zhang, Lei Xu, and Baowen Xu. RegCPython: a register-based Python interpreter for better performance. ACM Transactions on Architecture and Code Optimization, 20(1):14:1–14:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3568973.

Zhuo:2020:TEP

[ZZ20] Xilong Zhuo and Chenyi Zhang.

TFA: an efficient and precise virtual method call resolution for Java. Formal Aspects of Computing, 32(4– 6):395–416, November 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https://

link.springer.com/article/
10.1007/s00165-020-00518-
z.

Zhou:2023:RSA

[ZZdSO23] Litao Zhou, Yaoda Zhou, and Bruno C. d. S. Oliveira. Recursive subtyping for all. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):48:1–48:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3571241.